# Understanding the Local Geometry of Generative Model Manifolds

### v1

@@ Abstract
Deep generative models learn continuous representations of complex data manifolds using a finite number of samples during training. For a pre-trained generative model, the common way to evaluate the quality of the manifold representation learned, is by computing global metrics like Fréchet Inception Distance using a large number of generated and real samples. However, generative model performance is not uniform across the learned manifold, e.g., for foundation models like Stable Diffusion generation performance can vary significantly based on the conditioning or initial noise vector being denoised. In this paper we study the relationship between the local geometry of the learned manifold and downstream generation. Based on the theory of continuous piecewise-linear (CPWL) generators, we use three geometric descriptors – scaling ($\psi$), rank ($\nu$), and complexity ($\delta$) – to characterize a pre-trained generative model manifold locally. We provide quantitative and qualitative evidence showing that for a given latent, the local descriptors are correlated with generation aesthetics, artifacts, uncertainty, and even memorization. Finally we demonstrate that training a reward model on the local geometry can allow controlling the likelihood of a generated sample under the learned distribution.
@@
a beagle with a city
in the background
$\rho=0$                    $\rho=-1$                    $\rho=-1.5$
@@
Figure 1: Controlling sampling diversity using geometry guidance. We train a reward model on the geometric descriptor local scaling computed for the decoder of Stable Diffusion [22]. Positive (top-row) or negative guidance $\rho$ (bottom-row) via this reward model allows decreasing (top-row) or increasing (bottom-row) the likelihood of the generated samples under the learned distribution. As we decrease likelihood more background elements come into view and the focus on the subject decreases, vice-versa when increasing the likelihood.
@@
Figure 2: Exact geometry of a CPWL manifold. For a toy continuous piece-wise linear (CPWL) generator $\mathcal{G}:\mathbb{R}^{2}\rightarrow\mathbb{R}^{3}$, we provide analytically computed visualization of the input space partition, i.e., arrangement of linear regions (left) and learned CPWL manifold (middle-left). Each piece for this example, is colored by the piecewise-constant scaling induced by $\mathcal{G}$ that is analytically computed. Uniform samples from the latent domain (middle-right) and generated samples (right) are presented, colored by the estimated density at each sample using a gaussian kernel density estimator in $\mathbb{R}^{3}$. We see that for any sample $z\in\omega$, the estimated density ($\uparrow$ green) is inversely proportional with the scaling ($\downarrow$ green) for region $\omega$.
1 Introduction
In recent years, deep generative models have emerged as a powerful tool in machine learning, capable of synthesizing realistic data across diverse domains [15, 16, 22]. With the rapid increase in generation performance of such generative models, we have seen a demand for more fine-grained methods for evaluation. For example, global metrics of generation such as Fréchet Inception Distance (FID) are sensitive to both the fidelity and diversity of generated samples – precision and recall [23] were proposed to disentangle these two factors and quantify them separately. With the advent of foundational generative models like Stable Diffusion [22], human evaluation of images in a fine-grained manner has become the modus operandi for evaluation. In such cases, evaluating generative models extend beyond assessing mere sample fidelity and conditional alignment, to broader concerns about bias and memorization/duplication. The trend therefore shows a strong need for fine-grained understanding of the generative model behavior apart from what we can derive from global generation metrics. In this paper we look for answers to the following question:
Question.
For any sample generated using a pre-trained generative model, how is the local geometry of the learned manifold related to qualitative aspects of the sample?
We believe this question is especially significant 1) for models trained with large heterogeneous data distributions. Due to the heterogeneity, generation performance can vary significantly based on the sampling rate of the data manifold when training data is prepared. 2) For any pre-trained generative model without access to the training distribution, since computing global metrics like FID, precision or recall require the training data as ground truth.
How do we characterize the learned manifold?
To answer the question above, we require methods to derive local characteristics of a pre-trained generative model manifold for any given latent-sample pair $\{\bm{z},\bm{x}\}$. A large class of generative models – comprised of convolutions, skip-connections, pooling, or any continuous piece-wise linear (CPWL) operation – fall under the umbrella of CPWL generative models. Any CPWL operator, can be characterized exactly in-terms of the 1) knots/regions in the input space partition of the CPWL operator, where every knot denotes the location of a second-order change in the function and 2) the region-wise linear/affine operation, which scales, rotates, and translates any input region while mapping it to the output [1]. We therefore propose using the three geometric descriptors to characterize the learned manifold of a CPWL generator:
  • Local rank ($\nu$), that characterizes the local dimensionality of the learned manifold around $x$.
  • Local scaling ($\psi$), that characterizes the local change of volume by the generative model for an infinitesimal volume around $z$.
  • Local complexity ($\delta$), that approximates the smoothness of the generative model manifold in terms of second order changes in the input-output mapping.
Geometric descriptors such as local scaling, complexity or rank of Deep Neural Networks, have previously been used to measure function complexity [8] and expressivity [19, 20], to evaluate the quality of representations learned with a self-supervised objective [6], for interpretability and visualization of DNNs, [10], to understand the learning dynamics in reinforcement learning [4], to explain grokking, i.e., delayed generalization and robustness in classifiers [13], increase fairness in generative models [11], control sampling in generative models [12], and maximum likelihood inference in the latent space [18].
Our contributions.
In this paper, through rigorous experiments on large text-to-image latent diffusion models and smaller generative models, we demonstrate the correlation between the local geometric descriptors with generation aesthetics, diversity, and memorization of examples. We provide insights into how these manifest for different sub-populations under the generated distribution. We also show that the geometry of the data manifold is heavily influenced by the training data which enables applications in out-of-distribution detection and reward modeling to control the output distribution. Our empirical results present the following major observations, which can also be considered novel contributions of this paper:
  • C1. The local geometry on the generative model manifold is distinct from the off manifold geometry. Local scaling and local rank are indicative whether a $\{z,x\}$ pair is out-of-distribution.
  • C2. A strong relationship exists between the geometric descriptors and generation fidelity, aesthetics, diversity amd memorization in large text-to-image latent diffusion models.
  • C3. By training a surrogate model on the local geometry of Stable Diffusion, we can perform reward guidance to increase/decrease sampling uncertainty.
The paper is organized as follows: In Sec. 2 we define the local descriptors from first principles of CPWL generators. In Sec.3 we first show that for a toy DDPM [9] C1. holds. Following that we present visual and quantitative results showing C1. for Stable Diffusion. We also provide qualitative evidence for C2. for samples from the real Imagenet manifold. In Sec. 3.3.4 we draw connections between the local descriptors and 1) generation quality 2) uncertainty, and 3) memorization for Stable Diffusion. Finally in Sec. 4 we demonstrate how local geometry can be used as a reward signal to guide generation, providing evidence for C3.
2 Geometric Descriptors of the Learned Data Manifold
2.1 Continuous Piecewise-Linear Generative Models
Consider a generative network $\mathcal{G}$, which can be the decoder of a Variational Autoencoder (VAE) [17], the generator of a Generative Adversarial Network (GAN) [7] or an unrolled denoising diffusion implicit model (DDIM) [24]. Suppose, $\mathcal{G}:\mathbb{R}^{E}\rightarrow\mathbb{R}^{D}$ is a deep neural network with $L$ layers, input space dimensionality $E$ and output space dimensionality $D$. For any such generator, if the layers comprise affine operations such as convolutions, skip-connections, or max/avg-pooling, and the non-linearities are continuous piecewise-linear such as leaky-ReLU, ReLU, or periodic triangle, then the generator is a continuous piecewise-linear or piecewise-affine operator [1, 10]. This implies that the $\mathcal{G}:\mathbb{R}^{E}\rightarrow\mathbb{R}^{D}$ mapping can be expressed in terms of a subdivision of the input space into linear regions $\Omega$ and each region $\omega$ being mapped to the output data manifold via an affine opearation. The continuous data manifold or image of the generator $\text{Im}(\mathcal{G})$ can be written as the union of sets:
where, $\Omega$ is the partition of the latent space ${\mathbb{R}}^{E}$ into continuous piecewise-linear regions, $\bm{A}_{\omega}$ and $\bm{b}_{\omega}$ are the slope and offset parameters of the affine mapping from latent space vectors $z\in\omega$ to the data manifold. For the class of continuous piecewise-linear (CPWL) neural network based generative models, $\Omega$, $\bm{A}_{\omega}$, and $\bm{b}_{\omega}$ are functions of the neurons/parameters of the network. For a generator with $L$ layers, $\bm{A}_{\omega}$ and $\bm{b}_{\omega}$ can be expressed in closed-form in terms of the weights and the region-wise activation pattern of neurons for each layer. We refer the readers to Lemma 1 of [10] for details.
2.2 Local Geometry of Continuous Piecewise-Linear Generative Models
For any CPWL function, the input-output mapping can be expressed in terms of the input space subdivision $\Omega$ and the set of piecewise-affine parameters $\{\bm{A}_{\omega},\bm{b}_{\omega}\forall\omega\in\Omega\}$. In this section, we discuss how the local properties of the manifold can also be defined in terms of the regions $\omega$ or the region-wise affine parameters.
@@ 2.2.1 Local complexity, $\delta$
For CPWL functions, a general notion of complexity is based on the number of intervals or pieces of the function. For CPWL neural networks, a similar notion of complexity based on number of linear regions was proposed by [8]. We define local complexity of a CPWL generator as the following.

### v2

# What Secrets Do Your Manifolds Hold? Understanding the Local Geometry of Generative Models

@@ Abstract
Deep Generative Models are frequently used to learn continuous representations of complex data distributions using a finite number of samples. For any generative model, including pre-trained foundation models with Diffusion or Transformer architectures, generation performance can significantly vary across the learned data manifold. In this paper we study the local geometry of the learned manifold and its relationship to generation outcomes for a wide range of generative models, including DDPM, Diffusion Transformer (DiT), and Stable Diffusion 1.4. Building on the theory of continuous piecewise-linear (CPWL) generators, we characterize the local geometry in terms of three geometric descriptors - scaling ($\psi$), rank ($\nu$), and complexity/un-smoothness ($\delta$). We provide quantitative and qualitative evidence showing that for a given latent-image pair, the local descriptors are indicative of generation aesthetics, diversity, and memorization by the generative model. Finally, we demonstrate that by training a reward model on the local scaling for Stable Diffusion, we can self-improve both generation aesthetics and diversity using ‘geometry reward’ based guidance during denoising.
@@
a Australian terrier on
a cobblestone street
$\rho=0$                    $\rho=-1$                    $\rho=-3$
@@
Figure 1: Controlling visual complexity using geometry guidance. We train a reward model on the geometric descriptor local scaling computed for the decoder of Stable Diffusion (Rombach et al., 2021). Increasing local scaling, i.e., volume dilation (top-row) or increasing local volume contraction (bottom-row) of the denoised samples via guidance $\rho$, results in increased (top-row) or decreased (bottom-row) visual complexity of the generated samples. As we increase local scaling, more background elements come into view and the focus on the subject decreases, vice-versa when decreasing local scaling.
1 Introduction
In recent years, deep generative models have emerged as a powerful tool in machine learning, capable of synthesizing realistic data across diverse domains (Karras et al., 2019; 2020; Rombach et al., 2021). However, the performance of such generative models may not be uniform across all downstream tasks, for example, recent studies have demonstrated that models like Stable Diffusion can exhibit biases in generation in terms of reduced generation fidelity or diversity, for certain demographic groups Zhao et al. (2018); Luccioni et al. (2023). Especially for 1) models trained with large heterogeneous data distributions, and 2) any pre-trained generative model without access to the training distribution, the aforementioned observations can become hard to interpret or reason. Since generative models are essentially approximating a generative function mapping a latent space to the data manifold, one way to explain downstream performance could be in terms of the characteristics of the approximated function that can be measured using only the model internals. In this regard, we pose the following research question:
Research Question.
For any deep generative model pre-trained on an arbitrary dataset, how is the local geometry of the generator function related to downstream generation?
The theory of continuous piecewise-linear generators Balestriero et al. (2020) suggests that a large class of generative models can be considered continuous piecewise linear (CPWL) operators, implying that such generative models can be fully characterized in terms of their weights and architecture. We consider it the framework of choice to find answers to the aforementioned question and propose using three local geometric descriptors to quantify local characteristics of any pre-trained generative model:
  • Local rank ($\nu$), that characterizes the local dimensionality of the learned manifold.
  • Local scaling ($\psi$), that characterizes the local change of volume by the generative model input output mapping.
  • Local complexity ($\delta$), that approximates the un-smoothness of the generative model in terms of second order changes in the input-output mapping.
Geometric descriptors such as local scaling, complexity or rank, have previously been used to measure function complexity of Deep Neural Networks (DNN) (Hanin & Rolnick, 2019) and DNN expressivity (Poole et al., 2016; Raghu et al., 2017), to evaluate the quality of representations learned with a self-supervised objective (Garrido et al., 2023), for interpretability and visualization of DNNs (Humayun et al., 2023), to understand the learning dynamics in reinforcement learning (Cohan et al., 2022), to explain grokking, i.e., delayed generalization and robustness in classifiers (Humayun et al., 2024), maximum entropy or controlled sampling of GAN based generative models (Humayun et al., 2021; 2022b), and maximum likelihood inference in the latent space (Kuhnel et al., 2018).
Our contributions.
In this paper, through rigorous experiments on large image generative models, we establish correlations between local geometric descriptors and the aesthetic qualities, diversity, and degree of memorization of the generated samples. We demonstrate how these manifest differently for different sub-populations of the generative distribution. We also show that the geometry of the data manifold is heavily influenced by the training data which enables applications in out-of-distribution detection and reward modeling to control the output distribution. Our empirical results lead to the following conclusions:
  • C1. We present the first large-scale analysis of the local geometry of foundational text-to-image latent diffusion models and establish correlations between the local geometric descriptors and downstream aesthetic quality, diversity, and memorization (Sec 4.).
  • C2. For small diffusion models and foundational image generative models, we show that the local geometry on the generative model manifold is distinct from the off manifold geometry, and can help distinguish the domain of a generative model (Sec. 3).
  • C3. By training a auxiliary model on the local geometry of Stable Diffusion, we present a novel framework for reward guidance on a diffusion model to increase/decrease sampling diversity or control aesthetic qualities (Sec 5).
2 Local Descriptors of Generative Model Manifolds
We start by introducing the geometric descriptors we will use in our study and provide insight into what aspect of the generative model manifold geometry each of the descriptors quantify.
2.1 Continuous Piecewise-Linear Generative Models
Consider a generative network $\mathcal{G}$, which can be the decoder of a Variational Autoencoder (VAE) (Kingma & Welling, 2013), the generator of a Generative Adversarial Network (GAN) (Goodfellow et al., 2014), or an unrolled denoising diffusion implicit model (DDIM) (Song et al., 2020). Suppose, $\mathcal{G}:\mathbb{R}^{E}\rightarrow\mathbb{R}^{D}$ is a deep neural network with $L$ layers, input space dimensionality $E$ and output space dimensionality $D$. For any such generator, if the layers comprise affine operations such as convolutions, skip-connections, or max/avg-pooling, and the non-linearities are continuous piecewise-linear (CPWL) such as leaky-ReLU Xu (2015), ReLU, or periodic triangle, then the generator is a continuous piecewise-linear operator (Balestriero & Baraniuk, 2018a; Humayun et al., 2023). This implies that the $\mathcal{G}:\mathbb{R}^{E}\rightarrow\mathbb{R}^{D}$ mapping can be expressed in terms of a subdivision of the input space into linear regions $\Omega$ with each region $\omega$ from the input domain being mapped to the output via an affine operation. The continuous data manifold or image of the generator $\text{Im}(\mathcal{G})$ can be written as the union of sets:
where, $\Omega$ is the partition of the latent space ${\mathbb{R}}^{E}$ into continuous piecewise-linear regions, $\bm{A}_{\omega}$ and $\bm{b}_{\omega}$ are the slope and offset parameters of the affine mapping from latent space vectors $z\in\omega$ to the data manifold. For the class of continuous piecewise-linear (CPWL) neural network based generative models, $\Omega$, $\bm{A}_{\omega}$, and $\bm{b}_{\omega}$ are functions of the neurons/parameters of the network. For a generator with $L$ layers, $\bm{A}_{\omega}$ and $\bm{b}_{\omega}$ can be expressed in closed-form in terms of the weights and the region-wise activation pattern of neurons for each layer. We refer the readers to Lemma 1 of (Humayun et al., 2023) for details.
To help build intuition, without loss of generality lets consider a CPWL toy generator that is trained on a handcrafted task where the target function $f:\mathbb{R}^{2}\rightarrow\mathbb{R}^{3}$ is a mapping between $\mathbb{R}^{2}$ and a mixture of five gaussian functions. Since the learned function is a continuous piecewise-affine spline operator, we use SplineCAM (Humayun et al., 2023) to analytically compute the function learned by the generator and visualize the learned manifold, as well as the input space piecewise-linear partition learned by the generator in Fig. 2 middle-left and left. Each convex region $\omega$ bounded by the black lines, is mapped to $\text{Im}(\mathcal{G})$ via per region parameters as described in Equation 1. The input-output mapping operation by the generator is affine region-wise, therefore any given input space region can be scaled, rotated or translated with a continuity constraint between regions, while going from the input to the output. For CPWL generators there are three characteristics of the learned manifold that can be studied: i) the affine scaling induced per region, ii) the number of dimensions that are retained after scaling, i.e., local dimensionality of the learned manifold, and iii) the local smoothness of the CPWL partition. We now introduce local descriptors that can be used to characterize these quantities.
2.1.1 Local scaling, $\psi$
We first introduce local scaling as a target descriptor to be used in our study that measures the local scaling performed on a region $\omega$ by a CPWL generator.
Definition 1.
For a CPWL manifold produced by generator $\mathcal{G}$, the local scaling $\psi_{\omega}$ is constant within each region $\omega$, and measures the log-scaling of the volume induced by the affine slope $\bm{A}_{\omega}$ for all latents $\bm{z}\in\omega$. Local scaling for $\omega$ is expressed as
Refering back to the example in Fig. 2, each region on the CPWL manifold (middle-left) and in the input space (left) is colored by $\psi_{\omega}$, with darker shades indicating higher $\psi_{\omega}$. Suppose $\mathcal{G}$ has a uniform latent distribution, meaning every region $\omega$ has a uniform probability density in the latent space. Under an injectivity assumption for any input space region $\omega$ and $S=\{\bm{A}_{\omega}\bm{z}+\bm{b}_{\omega}\forall\bm{z}\in\omega\}$, according to Theorem 1. in Humayun et al. (2022a), the output density on $S$, $p_{S}(\bm{x})\propto\frac{1}{e^{\psi_{\omega}}}$. Therefore, local scaling $\psi_{\omega}$ is proportional to the negative log-likelihood of the generative model for any $z\in\omega$. We can validate this by using a kernel density estimator (KDE) to estimate the density of generated samples on the data manifold from a uniform latent distribution. In Fig. 2-right and middle-right, we denote the KDE estimated density per sample via colors, where higher density corresponds to brighter shades. Local scaling for any two regions $\omega\in\Omega$ and $\omega^{\prime}\in\Omega$, can therefore be used to express the difference in generation uncertainty for two locations on the data manifold:
$H_{\omega}-H_{\omega^{\prime}}=\psi_{\omega}-\psi_{\omega^{\prime}},$ (3)
2.1.2 Local rank, $\nu$.
The second descriptor we study is the rank of the region-wise slope matrix $\bm{A}_{\omega}$, which represents the dimensionality of the manifold learned by a CPWL generator.
Definition 2. For a CPWL manifold produced by generator $\mathcal{G}$, local rank $\nu_{\omega}$ is the exponent of the Shannon entropy of the spectral distribution of the per-region affine slope $\bm{A}_{\omega}$ and can be expressed as:
$\nu_{\omega}$ $=\exp\left(-\sum_{i}^{k}\alpha_{i}\log(\alpha_{i})\right)$ (4)
where $\text{where }\alpha_{i}$ $=\frac{\sigma_{i}}{\sum_{i}^{k}\sigma_{i}}+\epsilon.$ (5)
Here, $\{\sigma_{i}\}_{i=0}^{i=k}$ are non-zero singular values of $\bm{A}_{\omega}$ and $\epsilon=10^{-30}$ is a constant. The local rank $\nu_{\omega}$ can be shown to be equivalent to the dimensionality of the tangent space on the data manifold at $\bm{z}$.
@@ 2.1.1 Local scaling, $\psi$
Figure 2: The geometry of a continuous piecewise-linear toy generator. For a CPWL generator $\mathcal{G}:\mathbb{R}^{2}\rightarrow\mathbb{R}^{3}$, we provide analytically computed visualization of the input space partition, i.e., arrangement of linear regions (left) and learned CPWL manifold (middle-left). Each piece for this example, is colored by the piecewise-constant scaling induced by $\mathcal{G}$ that is also analytically computed. Uniform samples from the latent domain (middle-right) and generated samples (right) are presented, colored by the estimated density at each sample using a gaussian kernel density estimator in $\mathbb{R}^{3}$. We see that for any sample $z\in\omega$, the estimated density ($\uparrow$ green) is inversely proportional to the scaling ($\downarrow$ green) for region $\omega$.
@@ 2.1.3 Local complexity, $\delta$
An important geometric notion to characterize any manifold locally is the local smoothness of the manifold. However, smoothness requires computing the hessian of the input-output mapping making it computationally intractable for large generative models. We therefore consider local complexity as a proxy for sharpness of the manifold locally for our study. Based on the notion of complexity for CPWL neural networks (Hanin & Rolnick, 2019), we can define local complexity of a CPWL generator as the following.